A typical standard for military hikes is about 4 miles-per-hour with 30-inch strides per foot.

A) Estimate the frequency of foot strikes.

B) The definition of walking as opposed to running is that there is always contact with the ground. Estimate the duty cycle (fraction of time active) for the gastrocnemius muscle (calf) during the hike.

C) Based on your answer to A and B, what do you suppose the action potential train to the gastrocnemius muscle looks like?
Answer #1
  1. One step = 30 inches

Total distance covered = 4 miles =253440 inches/hr(1 mile = 63380inches)

Therefore total steps covered = 253440/30 = 8448 steps.

Frequency or f = no. of steps / second.

= 8448/3600 = 2.346 steps/ second will be the frequency of foot strikes

  1. Duty cycle = inverse of stride time

frequency = no. of steps per second

therefore we can find the time for one stride

=1/2.346

=0.426 sec

c. Result A and B shows that the military man are marching vigorously, that means there are 2.346 strides per second and it will cause faster contraction and relaxation of gastrocnemius muscle. There will be excitation-contraction coupling.
Answer #2

Bullshit 

That is wrong. EC coupling occurs whenever you are contracting muscles, not only when doing vigorous excercise. Its like saying theres an action potential because ofthe rate of walking lol. Part A was also wrong. The length of a step is twice that of a stride, so you need to multiply by 2. That was your most correct answer, and still wrong. 

Duty cycle is NOT the inverse of frequency. I dont know where you got that from. Duty cycle in muscles increases with frequency. 

https://www.pdf-archive.com/2017/08/30/joseph-j-feher-quantitative-human-physiology-2012/joseph-j-feher-quantitative-human-physiology-2012.pdf

Pg 270 

source: My cranial muscle
answered by: Bubinga Jones